# The Value of AI Advice: Personalized and Value-Maximizing AI Advisors Are Necessary to Reliably Benefit Experts and Organizations

### v1

@@ 1 Introduction
TeamRules: An Inherently Interpretable ReV-AI Advisor.
The ReV-AI learning framework can be used to derive different concrete methods for generating ReV-AI advisors of different model classes. In this work, we develop and evaluate a rule-based prototype algorithm, TeamRules (tr), which produces ReV-AI advisors that generate decision recommendations in the form of inherently interpretable rules. The tr advising process during deployment is shown in Figure 1b. Inherently interpretable models are especially suitable for and are commonly used to advise humans in high-stakes contexts [31, 33, 34, 35, 36, 37, 38, 39] because they faithfully represent the model’s actual decision-making process [40].
2 Results
Table 1: Summary of advisor generating methods, human behavior simulations, and measures of performance
Accounts for:
Advising
Framework
Advises
Selectively?
Human
Expertise?
Human
ADB?
Cost-benefit
Trade-off?
TR
TR-no(ADB)
TR-no(Cost)
TR-no(ADB, Cost)
Task-Only (Current Practice)
(a) Advisor Generators
Simulated Expertise Behaviors Simulated Confidence Behaviors Difficulty-Biased Decisions Group-Biased Decisions Accuracy-Biased ADB Group-Biased ADB Likelihood of human error varies with task difficulty [41]. Likelihood of human error varies across groups in the population [42, 43]. Human’s confidence in their independent judgment varies with but does not perfectly align with their decision accuracy [44, 45, 28]. Human confidence in their independent judgment varies across groups in the population [46].
(b) Types of Simulated Human Expertise and Confidence Behaviors
Metric Definition Value Added (Accuracy Improvement w.r.t. Human) -- (Cost-Benefit Trade-Off)*(Advising Costs Incurred) Accuracy Improvement w.r.t. Human (Number of Human Decisions Corrected Through Advising In Group) $\div$ (Total Number of Decision Instances) Advising Costs Incurred (Number of Pieces of Advice In Group) $\div$ (Total Number of Decision Instances) Advising Costs Incurred (In Accuracy Units [AU]) (Cost-Benefit Trade-Off)*(Advising Costs Incurred) Advising Confidence  [^2] Note that the confidence of rule-based advice corresponds to the precision of the rule. Thus, while the confidence is a measure of the quality of the rules used for advising and estimated prior to deployment, advising accuracy is a measure of the quality of decision advice offered under selective advising. tr selectively offers a rule as advice only if the advice is expected to add value. Hence, rules with any given precision are invoked selectively based on their expected added value, and may yield higher or lower advising accuracy. (Mean Confidence of Contradicting Advice for Instances Advised On In Group) Advising Accuracy (Number of Correct Pieces of Advice In Group) $\div$ (Total Number of Pieces of Advice In Group) Advising Rate (Number of Pieces of Advice In Group) $\div$ (Total Number of Decision Instances In Group) Advice Acceptance Rate (Number of Pieces of Advice Accepted By DM In Group) $\div$ (Number of Pieces of Advice Offered In Group)
(c) Measures of Performance
We conduct experiments to answer a set of key questions. First, we examine the value added by tr and alternative advisors to assess whether any of the them can be relied on to consistently add value across contexts. We also study what, if any, are the risks of ignoring the pillars, and what distinct advising strategies do tr’s advisors produce to reliably add value across contexts. Answering these questions requires consideration of a diverse set of contexts with different kinds of AI advisors produced to advise humans with varied human expertise and discretion behaviors in contexts with different advising cost-benefit trade-offs.
Alternative Advising Frameworks. We evaluated the value added by tr and other rule-based advisors that do not account for subsets of key pillars, including evaluation of the task-only advisor, which follows the current common practice in which AI models are trained to excel at a task by maximizing generalization accuracy and advising the human on all decisions. All advisor-generating frameworks considered, along with which pillars they account for, are shown in Table 1(c). The evaluations aim to reveal the potential performance benefits and risks of alternatives that do not consider all or some of the key pillars.
Figure 2: Value added by different AI advisors relative to independent human for environments with varying trade-offs between advising costs and decision benefits for Heart Disease, FICO, and HR datasets. Value added is the improvement in human’s decisions from advising minus costs incurred from human engagements. X-axis reflects contexts in which the cost from 1/x human engagements are equivalent to the loss from one incorrect decision: higher x-values reflect that fewer human engagements are permissible to achieve a given decision benefit. task-only model generalization accuracy is 0.78 (Heart Disease), 0.71 (FICO), 0.79 (HR). The human’s standalone generalization accuracy for each dataset and decision behavior is 0.71 (Heart Disease Difficulty-Biased), 0.83 (Heart Disease Group-Biased), 0.73 (FICO Difficulty-Biased), 0.70 (FICO Group-Biased), 0.70 (HR Difficulty-Biased), 0.71 (HR Group-Biased). Results show average value-added +- SE (shaded region) over 10 repetitions.
Data. We use three data domains which in practice represent consequential decisions: the Heart Disease Dataset in which labels reflect the presence of heart disease for a patient [47], the FICO dataset [48] reflecting loan applications to inform loan decisions, and an HR employee attrition dataset in which labels reflect employee departure, and thus such data can be used in practice to inform employment retention decisions. [^3] https://www.kaggle.com/datasets/pavansubhasht/ibm-hr-analytics-attrition-dataset, note this is a fictional dataset that represents a real-world task
Simulations. To explore AI advising performance for a wide range of settings, and similarly to recent work that developed methods for human-AI collaboration [49, 32, 50, 51, 52, 53], we simulate diverse human expertise [54, 49, 55, 53], confidence, and corresponding ADB [44, 45, 28, 46] informed by prior work on human behavior, summarized in Table 1(c). Complete details of the simulation procedures and parameters are outlined in Appendix A. We build on prior work that modeled and validated human ADB with real human interactions [27], and which maps the human’s self-reported confidence and the AI advisor’s confidence to the human’s likelihood to accept the AI’s advice. In general, humans are more likely to accept AI advice the higher the AI’s confidence is relative to their own confidence in their decision. We simulate various human confidence patterns from previous studies [44, 45, 28], which lead to different probabilities of accepting AI advice based on the AI’s confidence level for each instance.
Evaluations. For different contexts characterized by different data domains, human expertise and ADB, Figure 2 shows the AI advisors’ added-value relative to the human’s as a function of the context’s trade-off: the cost incurred from reconciling one disagreement is equivalent to the loss from $x$ decision errors. Thus, higher values on the x-axis reflect contexts in which fewer human engagements are tolerated to achieve a given decision benefit, where a human engagement refers to the human reconciling AI advice for a single instance, and AI advice refers to an AI decision recommendation that contradicts the human’s initial judgment. The dashed horizontal line is the value from the human’s independent decisions. All evaluation metrics are summarized in Table 1(c) and discussed in detail in Appendix A.4.
The value-added and lost by AI advisors.
As demonstrated across diverse contexts in Figure 2, tr’s advisors consistently add the most value and, where not superior, perform comparably to alternatives. [^4] In Appendix C.1 and C.2, we provide results for additional settings with even more inhibitive advising cost-benefit trade-offs and with different decision loss cost structures. tr advisors can also be relied on to not diminish the value offered by the human alone. Importantly, we see that tr’s performance is consistent even in contexts where all other AI advisors struggle to add value. For example, it is more difficult to add value in contexts where fewer human engagements are tolerated to achieve a given improvement in decision outcomes (higher x-values), reflecting contexts such as emergency medicine. Yet, in such contexts, tr advisors add the most value and avoid losses from cost-ineffective human engagement. It is also challenging for AI advisors to add value in typical expert settings where the expert’s strong decision-making abilities and confidence in their own judgment limit opportunities for improvement, given this entails producing even more accurate and highly persuasive advice. For example, in Figure 2b, the task-only AI advisor always diminishes value, even when the context incurs no costs from human engagement (x-axis value of 00), demonstrating the model class’s difficulty of learning advice that is superior and convincing enough to improve the human’s decisions at all, let alone cost-effectively when advising costs are less tolerated. To benefit such contexts it is especially crucial to learn and selectively offer value-maximizing advice that is simultaneously sufficiently accurate and convincing to warrant the human’s attention, as demonstrated by tr.
Figure 3: Case Study - Outcomes: Advising outcomes when tr, task-only, tr-no(adb) advisors advise an over-confident expert. The context’s trade-off is $0.1$, reflecting that reconciling up to 10 contradictory pieces of AI advice is deemed cost-effective if the contradictions yield at least one improved decision. Advising costs incurred shown are converted to units of decision loss given the trade-off (AU). The expert has an independent decision accuracy of 90% on the Male population (majority), a 60% accuracy on the Female population (minority), and an overall accuracy of 87.5% on the entire population. This expert is highly confident (97.5% confidence) on the entire population. Results are averages over 20 repetitions $\pm$ SE (vertical line at center of each bar).
Figure 4: Case Study - Advising Strategies: Properties of advice corresponding to Figure 3. Advising overconfident expert leads TR to learn and selectively offer higher accuracy and more confident advice that the expert is more likely to accept, leading to more improvements in decisions from fewer expert engagements. Results from 20 repetitions.
Our results also reveal that AI advisors that do not advise selectively or are not informed by one or more of the pillars not only fail to maximize their value but also risk diminishing the value of the context in which they operate. Task-only advisors, which are commonly deployed in practice, incur the greatest losses, even when their generalization accuracy exceeds the human’s independent accuracy, highlighting the risks associated with the most prevalent AI advisors, and consistent with evidence from fields like radiology, where increased human engagement to reconcile disagreements has often been deemed cost-ineffective [5]. Other advisors, such as tr-no(adb) which includes all but one of tr’s pillars and consequently does not attempt to learn advice that is sufficiently inherently convincing to the human, can also diminish the value offered by the human alone when the context is not tolerant of unlimited human engagements (high x-axis values in Figures 2b,d,e,f,g,h).
We find that when neglecting the human’s ADB, tr-no(adb) matches tr’s value when unlimited human engagements are allowed to achieve a given decision improvement (x-axis value of 0). However, as the context permits fewer engagements to achieve a given decision benefit (higher x-axis values), accounting for the human’s ADB becomes essential to add value. This is because, to be cost-effective, advice must not only be superior but also sufficiently convincing to justify the cost of engagement. Accounting for the human’s ADB is essential to accurately estimate the likelihood of the advice being convincing and accepted. We also find that accounting for the human’s decision-making behavior can be beneficial even in contexts allowing unlimited human engagement. This is evidenced by the relative value added by tr-no(adb,cost) (which considers the human’s expertise) compared to task-only. The advantage of tr-no(adb,cost) stems from its ability to focus on decisions the human is less likely to handle correctly and to withhold advice when it is less likely to be accurate than the human’s judgment. Importantly, our findings reveal that AI advisors which do not account for the pillars can diminish value—even when using models that exceed human accuracy, while, interestingly, ReV-AI advisors can reliably add value to their contexts, even when the preferred model class cannot surpass human accuracy. Overall, the results demonstrate that the system-level, value-based approach we advocate for is essential to produce AI advisors that are dependable and can reliably benefit contexts, while highlighting the risks associated with alternative approaches.
How TeamRules’ distinct advising strategies benefit experts?
We now examine the distinct advising strategies that tr learns to reliably add value compared to alternative approaches and share insights on the role of the pillars and the selective advising mechanism. Towards this, we consider the Heart Disease diagnosis (classification) task [47], and a context inspired by evidence of physicians misdiagnosing heart failure in women at a higher rate as compared to men [56, 43, 57]. We simulate an expert with the reported performance bias who is also highly confident in their diagnoses for both men and women. While the expert would benefit from cost-effectively improving their decisions on women, it is challenging to produce advice that is inherently convincing to the expert given it ought to be even more confident than the expert’s (miscalibrated) overconfidence. The expert’s overconfidence thus limits the scope to improve their decisions cost-effectively, yet, such improvements would be consequential.
Figure 4 shows measures reflecting the advising strategies produced by the AI advisors and Figure 3 shows the outcomes when each AI advisor is paired with the expert to produce decisions. To characterize tr’s distinct advising strategy we contrast it with that of the task-only and tr-no(adb) advisors. We find that tr adds the most value by both improving more of the expert’s final decisions (Figure 3d,e,f) while achieving this through fewer engagements (Figure 3g,h,i). tr’s distinct advising strategy is shaped by each of the pillars, including the expert’s overconfident ADB, decision behavior, and context’s trade-off. Specifically, note that given the expert’s near perfect decision accuracy and high confidence for male patients, the scope for cost-effective advising on men is rather limited. We find that tr advises the expert highly selectively relative to the alternatives (Figure 4g,h,i). While most of the patients are males, tr’s advice focuses on women (65% of the advice), which is the group for which there are more opportunities to benefit the context and for which tr’s advice ultimately adds the most value (Figure 3a). As shown, tr’s conservative advising strategy on men yields a small improvement in decisions for men while avoiding the large losses from cost-ineffective human engagements incurred by the alternatives.
We also learn how the nature of tr’s advice is distinctly advantageous. tr advises with highly accurate (Figures 4d,e,f) and confident advice relative to alternatives (Figures 4a,b,c), leading the overconfident expert to be around twice as likely to accept tr’s advice relative to task-only’s advice (Figures 4j,k,i), and thereby improves more decisions cost-effectively. In contrast, the task-only advisor engages the human much more often with contradictory, yet, less accurate and less confident advice on both populations, ultimately harming value. Additionally, the tr-no(adb) advisor, which learns and selectively offers advice tailored to the human’s expertise and the context’s trade-off, and thereby provides highly accurate advice focusing on the female population, does not provide advice that is simultaneously sufficiently inherently convincing to the expert. Consequently, tr-no(adb)’s advice is accepted less often, and is thereby not cost-effective enough to warrant the human’s engagement.
In Appendix C.4, we discuss a complementary case study where the expert’s lower confidence makes them more receptive of advice for female patients. In this case, tr adapts to the human’s receptiveness by advising the expert on a higher proportion of female patients with lower-confidence advice that is sufficiently accurate and persuasive to cost-effectively improve more of the expert’s decisions..
3 Discussion
Although AI advisors can outperform human experts in some tasks, they risk diminishing value across AIaDM contexts. This can harm high-stakes practices, erode trust in AI, and slow progress. To enhance the positive impact of AI advisors, we advocate for a system-level, value-driven perspective. This shift is essential to understanding the failures observed in practical applications and to developing AI advisors that are reliable. Central to this approach are key pillars that characterize how AI advice can either add to or detract from value in practice. Our results establish that Rev-AI advisors can be relied on to add value across settings, and that neglecting any pillar risks diminishing value. The pillars also highlight the challenges of creating value-adding systems, such as understanding the human engagement costs, aligning with the context’s tolerance for those costs, and recognizing how the inherent properties of advice, besides its accuracy, influence its acceptance and added value.
This work thus provides critical insight into why AI advising systems struggle to add value in practice, and underscores that a system-level and value-based perspective is crucial to reliably add value in AIaDM contexts and enhance trust in AI systems. Our work provides timely and actionable guidance and a value-based framework to catalyze research on AI advising systems and foster their positive impact in practice.
4 Methods
The ReV-AI learning framework can be used to derive different concrete methods for generating ReV-AI advisors of different model classes. In this work, we develop and evaluate a rule-based prototype algorithm, TeamRules (tr), which produces ReV-AI advisors that generate decision recommendations in the form of inherently interpretable rules. The tr advising process during deployment is shown in Figure 1b. Inherently interpretable models are especially suitable for and are commonly used to advise humans in high-stakes contexts [31, 33, 34, 35, 36, 37, 38, 39] because they faithfully represent the model’s actual decision-making process [40].
@@ 4.1 Problem Formulation
We now formalize the framework for producing ReV-AI advisors shown in Figure 1a. Key notation for this section is summarized in Table 2. Let $\mathcal{X}$ be the set of decision instances with labels $\mathcal{Y}=\{0,1\}$, where $\mathcal{X}\times\mathcal{Y}\sim{\mathcal{D}}$. A decision instance $i$ is represented by features $x_{i}$, and outcome $y_{i}$. We consider contexts reflecting common high-risk decision domains in which a human first makes an independent decision $h_{i}\in\{0,1\}$. The human is then advised by an AI advising model, denoted by $m$, that either withholds advice or offers advice in the form of a decision recommendation $\hat{y}_{i}$. The AI recommendation is complemented by the AI model’s confidence [^5] We consider the model’s confidence because the impact of model confidence on human likelihood to accept advice has been studied in prior work. The framework can be instantiated with other forms of complementary ADB-bound information under the condition that the impact the ADB-Bound information has on ADB is modeled. $c^{M}_{i}$ reflecting the AI’s assessment of the likelihood that it’s advice is accurate. Finally, the human produces a final decision, by either accepting the recommendation $\hat{y}_{i}$ or selecting their own independent decision $h_{i}$.
@@
$\hat{p}(a|c^{M},c^{H},h\neq\hat{y}))$, denoted with $\hat{p}(a)$ for brevity Estimated model of human’s algorithm discretion behavior (ADB) - likelihood human accepts a contradicting recommendation given their confidence in their initial decision and model’s confidence in its advice.
@@
$\mathcal{L}(y,\hat{y},h,\hat{p}(a),\alpha)$ Expected team loss over human’s likelihood to accept a piece of advice.
@@
$\mathbbm{C}(R,x_{i})$ Rule coverage indicator. True if instance with feature values $x_{i}$ is covered by any rule in $R$.
@@
$\mathbbm{1}\{\text{condition}\}$ Indicator function. Returns $1$ if condition is met, 00 otherwise.
@@
ReV-AI’s advice $\hat{y}$ for a decision instance is either a decision recommendation $z\in\mathcal{Y}$ or the lack of a decision recommendation, such that the withholding of advice corresponds to recommending the human’s initial judgment $h$. Formally, $\hat{y}$ is given by:
$\hat{y}$ =$=\begin{cases}z\in\mathcal{Y},&\text{if decision advice is offered}\\ h,&\text{if advice is withheld}\end{cases}$ (1)
Simultaneously, the selective advising decision – whether to provide advice and the specific advice $z$ – is personalized to the human and context’s cost-benefit trade-off, with the goal of maximizing the advisor’s added value:
@@
where $\mathbb{E}_{x,y,h\sim\mathcal{D}}$ denotes expectation over the input space $(X,Y)$ and the human’s independent decisions over it with joint distribution $D$; $\mathcal{L}$ is the expected team loss over the human’s discretion $a$ resulting from the given AI advice, the given instance $(x,y)$, and and the human’s independent decision $h$, given by:
$\mathcal{L}(y,\hat{y},h,\hat{p}(a),\alpha)$ $=\hat{p}(a)\mathcal{V}\big{(}y,\hat{y}\big{)}$ (3)
$\text{\indent}+\big{(}1-\hat{p}(a)\big{)}\mathcal{V}(y,h)+\alpha\mathbbm{1}\{\hat{y}\neq h\}$
@@
where $\mathcal{V}(y,d)$ denotes the context-relevant loss incurred from the human’s final decision $d$ given ground truth $y$, and $\alpha$ denotes the context’s trade-off between reconciling contradictory advice and benefits from improving decisions, such that the loss from reconciling one contradictory advice is equivalent to the following loss from a decision error: $\frac{\alpha}{\mathcal{V}(y,d|y\neq d)}$. Equivalently, the trade-off implies that $\frac{\mathcal{V}(y,d|y\neq d)}{\alpha}$ contradicting pieces of advice produce loss equivalent to that of a single incorrect decision. Note that Equation (3) reflects that when $m$ offers contradictory advice, i.e., when $\hat{y}\neq h$, the expected loss depends on the estimated likelihood $\hat{p}(a)$ that the human accepts the advice; the loss in this case also includes the cost $\alpha$ from the human’s reconciliation of the contradiction. When $m$ does not provide decision advice or when it provides advice that agrees with the human’s initial judgment $h$, then $\hat{y}=h$, and the loss becomes $\mathcal{V}(y,h)$.
@@
The total expected loss produced by an advisor, $\mathbb{E}_{x,y,h\sim\mathcal{D}}\bigg{(}\mathcal{L}(y,\hat{y},h,\hat{p}(a),\alpha)\bigg{)}$, is henceforth referred to as the Total Team Loss (TTL). Importantly, as reflected in our discussion and shown in Equation (3), the loss includes the loss from reconciling contradictions in addition to the loss from incorrect final decisions. Ultimately, offering value-adding advice hinges on improvements achieved cost-effectively per the context’s cost-benefit trade-off.
@@
The ReV-AI advisor is informed by the underlying data domain $\mathcal{X}\times\mathcal{Y}$, data reflecting the human’s expertise, confidence, and ADB for estimating $p(a|c^{M},c^{H},\hat{y}\neq h)$, and the context’s tolerance for added costs to improve decisions. Data reflecting the human’s expertise - consisting of the human’s independent decisions =$\mathcal{H}=\{0,1\}$ for instances in $\mathcal{X}$ is often available in historical data and which otherwise can be acquired for the corresponding domain. Prior work also established how a given individual’s likelihood of accepting a given piece of AI advice for a given decision, $\hat{p}(a)$, can be effectively estimated based on the human’s self-reported (possibly miscalibrated) confidence in their decision and the AI’s confidence, $c^{H}$ and $c^{M}$, respectively [27, 28]. Following prior work, for each instance $(x,y)$ for which advice is offered, the human’s decision $h$, self-reported confidence $c^{H}$, and decision to accept or reject the advice $a$, is acquired in a risk-free setting, prior to deployment. Prior to producing an initial ReV-AI advisor, data can be acquired for advice offered by a general-purpose model trained on $\mathcal{X}\times\mathcal{Y}$, based on which $\hat{p}(a)$ can be modeled and an initial ReV-AI can be produced. Importantly, data on both human expertise and discretion can be acquired periodically to reflect changes. We discuss modeling the human’s discretion in more detail in Appendix A.2 and we study the robustness of the approach to varying degrees of accuracy for these predictions in Appendix C.5.

### v2

@@ 2.4 Problem Formulation
We now formalize the framework for producing ReV-AI advisors shown in Figure 2. Key notation for this section is summarized in Table 1. Let $\mathcal{X}$ be the set of decision instances with labels $\mathcal{Y}=\{0,1\}$, where $\mathcal{X}\times\mathcal{Y}\sim{\mathcal{D}}$. A decision instance at index $i$ is represented by features $x_{i}$, and outcome $y_{i}$. We consider contexts reflecting common high-risk decision domains in which a human first makes an independent decision $h_{i}\in\{0,1\}$. The human is then advised by an AI advising model, denoted by $m$, that either withholds advice or offers advice in the form of a decision recommendation $\hat{y}_{i}$. The AI recommendation is complemented by the AI model’s confidence [^3] We consider the model’s confidence because the impact of model confidence on human likelihood to accept advice has been studied in prior work. The framework can be instantiated with other forms of complementary ADB-bound information under the condition that the impact the ADB-Bound information has on ADB is modeled. $c^{M}_{i}$ reflecting the AI’s assessment of the likelihood that its advice is accurate. Finally, the human produces a final decision, by either accepting the recommendation $\hat{y}_{i}$ or selecting their own independent decision $h_{i}$.
@@
$)\hat{p}(a|c^{M},c^{H},h\neq\hat{y}$)), $\hat{p}(a)$ for brevity Estimated model of human’s algorithm discretion behavior (ADB) - likelihood human accepts a contradicting recommendation given their confidence in their initial decision and model’s confidence in its advice.
@@
$\mathcal{L}(y,\hat{y},h,\hat{p}(a),\alpha)$ Expected team loss from one AI advice over human’s likelihood to accept it
@@
$\mathbb{C}(R,x_{i})$ Rule coverage indicator. True if instance with feature values $x_{i}$ is covered by any rule in $R$.
@@
$\mathbb{I}\{\text{condition}\}$ Indicator function. Returns $1$ if condition is met, 0 otherwise.
TDL Team decision loss from final decisions (loss from decisions only, does not include advising costs).
@@
A ReV-AI advisor’s advice $\hat{y}$ for a decision instance is either a decision recommendation $z\in\mathcal{Y}$ or the lack of a decision recommendation, such that the withholding of advice corresponds to recommending the human’s initial judgment $h$. Formally, $\hat{y}$ is given by:
$\hat{y}$ $=\begin{cases}z\in\mathcal{Y},&\text{if decision advice is offered}\\ h,&\text{if advice is withheld}\end{cases}$ (1)
Simultaneously, the selective advising decision – whether to provide advice and the specific choice of advice to provide $z$ – is personalized to the human and context’s cost-benefit trade-off, with the goal of maximizing the advisor’s added value:
@@
where $\mathbb{E}_{x,y,h\sim\mathcal{D}}$ denotes expectation over the input space $(X,Y)$ and the human’s independent decisions over it with joint distribution $D$. $m^{*}$ is thus the model from a set of possible models $\mathcal{M}$ that minimizes $\mathbb{E}_{x,y,h\sim\mathcal{D}}(\mathcal{L}(y,\hat{y},h,\hat{p}(a),\alpha))$. $\mathcal{L}$ is the Expected Team Loss over the human’s discretion $a$ resulting from the AI advice $\hat{y}$, the decision instance $(x,y)$, and the human’s independent decision $h$, given by:
[…]
$\text{\indent}+\big(1-\hat{p}(a)\big)\mathcal{V}(y,h)+\alpha\mathbb{I}\{\hat{y}\neq h\},$
@@
where $\mathcal{V}(y,d)$ denotes the context-relevant loss incurred from the human’s final decision $d$ given ground truth $y$. $\mathcal{V}(y,d)$ can be selected to fit the context, and the framework makes no assumptions on the loss produced by different types of errors. $\alpha$ denotes the context’s trade-off between reconciling contradictory advice and benefits from improving decisions, such that the loss from reconciling one contradictory advice is equivalent to the following loss from a decision error: $\frac{\alpha}{\mathcal{V}(y,d|y\neq d)}$. Equivalently, the trade-off implies that $\frac{\mathcal{V}(y,d|y\neq d)}{\alpha}$ contradicting pieces of advice produce loss equivalent to that of a single incorrect decision. Note that Equation (3) reflects that when $m$ offers contradictory advice, i.e., when $\hat{y}\neq h$, the expected loss depends on the estimated likelihood $\hat{p}(a)$ that the human accepts the advice; the loss in this case also includes the cost $\alpha$ from the human’s reconciliation of the contradiction. When $m$ does not provide decision advice or when it provides advice that agrees with the human’s initial judgment $h$, then $\hat{y}=h$, and the loss becomes $\mathcal{V}(y,h)$.
@@
The total expected loss produced by an AI advisor, $\mathbb{E}_{x,y,h\sim\mathcal{D}}\bigg(\mathcal{L}(y,\hat{y},h,\hat{p}(a),\alpha)\bigg)$, is henceforth referred to as the Total Team Loss (TTL). Importantly, as reflected in our discussion and shown in Equation (3), the loss includes the loss from reconciling contradictions in addition to the loss from incorrect final decisions. Ultimately, offering value-adding advice hinges on improvements achieved cost-effectively per the context’s cost-benefit trade-off.
@@
The ReV-AI advisor is informed by the underlying data domain $\mathcal{X}\times\mathcal{Y}$, data reflecting the human’s expertise, confidence, and ADB for estimating $p(a|c^{M},c^{H},\hat{y}\neq h)$, and the context’s tolerance for added costs to improve decisions. Data reflecting the human’s expertise - consisting of the human’s independent decisions $\mathcal{H}=\{0,1\}$ for instances in $\mathcal{X}$ - is often available in historical data and which otherwise can be acquired for the corresponding domain. Prior work also established how an individual’s ADB – i.e., likelihood of accepting a given piece of AI advice for a given decision, $\hat{p}(a)$ – can be effectively estimated based on the human’s self-reported (possibly miscalibrated) confidence in their decision and the AI’s confidence, $c^{H}$ and $c^{M}$, respectively (will_you_accept, ; CHONG2022107018, ).
2.5 Data Acquisition, Training, and Evaluation
In most high-stakes contexts, regulation and best practices dictate that the efficacy and safety of new technologies should be evaluated prior to deployment to mitigate harms. We follow this paradigm and consider a flow in which the expert initially interacts with AI advice over time in a risk-free environment, prior to producing the AI advisor. During this phase, data is collected, after which a ReV-AI advisor is trained, and subsequently the efficacy and safety of the advisor are evaluated. Concretely, building on prior work (will_you_accept, ; CHONG2022107018, ), for a subset of risk-free training decision instances $(x,y)$, the human’s decision $h$, self-reported confidence $c^{H}$, and decision to accept or reject the AI advice $a$, are acquired and used to model the human’s ADB. AI advice is initially offered by a general-purpose model trained on the joint space $\mathcal{X}\times\mathcal{Y}$. The remaining set of risk-free training decision instances are used to train the ReV-AI advisor. Complete details of modeling the human’s ADB is outlined in Appendix A.2.
3 Methods: TeamRules, An Inherently Interpretable ReV-AI Advisor
The ReV-AI learning framework can be used to derive different concrete methods for generating ReV-AI advisors of different model classes. In this work, we develop and evaluate a rule-based prototype algorithm, TeamRules (tr), which produces ReV-AI advisors that generate decision recommendations in the form of inherently interpretable rules. Inherently interpretable models are especially suitable for and are commonly used to advise humans across high-stakes contexts (rudin2019stop, ; wang2021scalable, ; rich, ; BALAGOPAL2021102101, ; edit_machine, ; pmlr-v139-biggs21a, ; Subramanian_Sun_Drissi_Ettl_2022, ; prescriptive_relu, ) because they faithfully represent the model’s actual decision-making process (dangers_laugel, ). Rule-based models also present challenges for optimization, hence we also show how these challenges can be effectively addressed. The tr advising process during deployment is shown in Figure 3.
Figure 3: TeamRules Advising Process in Deployment
Independently, as done in practice with non-stationary concepts, such as human preferences or choices and market outcomes, changes in the underlying concept merely entail degradation in the model’s performance which can be monitored to trigger retraining and evaluation, as desirable. Importantly, our framework does not assume perfect estimates of ADB and we empirically demonstrate that producing effective ReV-AI advisors does not rely on perfect estimates. In the evaluations reported here we induce the discretion model from data and later study how increasing mis-estimates of the human’s ADB impact the AI advisor’s performance and benefits.